Differential Analysis Involving Opportunity Costs

On July 1, Coastal Distribution Company is considering leasing a building and buying the necessary equipment to operate a public warehouse. Alternatively, the company could use the funds to invest in $740,000 of 5% U.S. Treasury bonds that mature in 14 years. The bonds could be purchased at face value. The following data have been assembled:

Cost of store equipment $740,000
Life of store equipment 14 years
Estimated residual value of store equipment $75,000
Yearly costs to operate the warehouse, excluding
depreciation of store equipment $175,000
Yearly expected revenues—years 1-7 $280,000
Yearly expected revenues—years 8-14 $240,000

Required:

1. Prepare a differential analysis as of July 1 presenting the proposed operation of the warehouse for the 14 years (Alternative 1) as compared with investing in U.S. Treasury bonds (Alternative 2). If an amount is zero, enter "0". For those boxes in which you must enter subtracted or negative numbers use a minus sign.

Differential Analysis
Operate Warehouse (Alt. 1) or Invest in Bonds (Alt. 2)
July 1
Operate Warehouse (Alternative 1) Invest in Bonds (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues $ $ $
Costs:
Costs to operate warehouse
Cost of equipment less residual value
Income (Loss) $ $ $

Feedback

Subtract the warehouse costs (14 years) and the cost of the equipment less the residual value from the revenues from operating the warehouse. Determine the bond investment interest income for 14 years (principal × rate × time). Determine the differential effect on income of the revenues, costs, and income (loss) by subtracting alternative 2 from alternative 1. Which alternative has the most desirable effect on income?

Learning Objective 1.

2. Based on the results disclosed by the differential analysis, should the proposal to operate the warehouse be accepted?
Yes

3. If the proposal is accepted, what is the total estimated income from operations of the warehouse for the 14 years?
$
Answer #1
Operate Warehouse (Alt. 1) or Invest in Bonds (Alt. 2)
July 1
Operate Warehouse (Alternative 1) Invest in Bonds (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues $3,640,000 $518,000 $-3,122,000
Costs:
Costs to operate warehouse -$2,450,000 0 $2,450,000
Cost of equipment less residual value -$665,000 0 $665,000
Income (Loss) $525,000 $518,000 -$7000

1. revenues = (280000*7)+(240000*7)    7400000*5%*14years = 518000

=1960000+1680000

=$3,640,000

2.cost   = 175000*14=$2,450,000

3. 740000-75000= $665,000

requirement 2.

yes. it should be accepted as it results in more income as compare to investing in US bonds.

requirement 3.

estimated income from operating warehouse for 14 years is $525,000